# VersaTune: An Efficient Data Composition Framework for Training Multi-Capability LLMs

### v4

@@ Abstract.
Large-scale pretrained models, particularly Large Language Models (LLMs), have exhibited remarkable capabilities in handling multiple tasks across domains due to their emergent properties. These capabilities are further augmented during the Supervised Fine-Tuning (SFT) phase. Despite their potential, existing work mainly focuses on domain-specific enhancements during fine-tuning, the challenge of which lies in catastrophic forgetting of knowledge across other domains. In this study, we introduce VersaTune, a novel data composition framework designed for enhancing LLMs’ overall multi-ability performances during training. We categorize knowledge into distinct domains including law, medicine, finance, science, code, etc. We begin with detecting the distribution of domain-specific knowledge within the base model, followed by the training data composition that aligns with the model’s existing knowledge distribution. During the training process, domain weights are dynamically adjusted based on their learnable potential and forgetting degree. Experimental results demonstrate that VersaTune achieves significant improvements in multi-domain performance, with an enhancement in comprehensive multi-domain tasks. Additionally, in scenarios where specific domain optimization is required, VersaTune reduces the degradation of performance in other domains by , while preserving the target domain’s training efficacy.
PVLDB Reference Format:
Keer Lu, Keshi Zhao, Zheng Liang, Da Pan, Shusen Zhang, Xin Wu, Zenan Zhou, Guosheng Dong, Bin Cui, Wentao Zhang. PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX
† Corresponding Author.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and other artifacts will be released after the paper is accepted.
1. Introduction
@@
In recent years, large-scale pretrained models with enormous pretraining tokens and parameters, such as Transformer models, have become a cornerstone in Artificial Intelligence (AI) (Achiam et al., 2023; Dwivedi et al., 2021; Lewkowycz et al., 2022), particularly for Natural Language Processing tasks (Brown et al., 2020; Devlin, 2018; Radford et al., 2019), reshaping AI research and applications in domains such as law (Cui et al., 2023), medicine (Singhal et al., 2023; Thirunavukarasu et al., 2023), finance (Li et al., 2023b; Wu et al., 2023), science (Beltagy et al., 2019; Taylor et al., 2022) and code (Liu et al., 2024; Roziere et al., 2023). The success of Transformer models, such Large Language Models (LLMs), stems from their capabilities to automatically learn and distill hierarchical data representations, making them highly effective for complex tasks (Nie et al., 2023). In order to further enhance such abilities across these areas, LLMs typically undergo the supervised fine-tuning (SFT) stages on domain-specific datasets.
Data plays a fundamental role throughout the training phases of LLMs. While the significance of model architectures is widely acknowledged, data-centric approaches hold substantial importance (Wang et al., 2023), which emphasizes the quality, organization, and management of data as the crucial factor for improving model performance (Miao et al., 2024; Fernandez et al., 2023; Whang et al., 2023). Efficient data management, especially in constructing an appropriately tailored training dataset with suitable data composition, is critical for synergizing the model’s performance and ability to generalize across different tasks and domains during both pretraining and supervised fine-tuning stages (Wang et al., 2023; Kumar et al., 2017).
@@
As demonstrated by the robust performances of state-of-the-art LLMs such as GPT-4 (Achiam et al., 2023), PaLM (Anil et al., 2023) and Gemini (Team et al., 2023), LLMs have the potential to master multiple tasks across all specific domains within a single model. However, most existing research on supervised fine-tuning tends to merely concentrate on a single ability of LLMs (Dong et al., 2023; Xu et al., 2024), with the multi-task performance on composite data of essentially different downstream tasks being less studied. We try to enhance the overall multitasking performance of LLMs across various domains by optimizing data mixing ratios during training:
@@
How to design a data composition strategy that could achieve overall multitasking capabilities?
@@
Through extensive experimental analysis, we identified that the challenges associated with data composition strategies stem from the following three key aspects:
@@
C1: Catastrophic Forgetting. Given the fundamental differences between tasks across various domains, for multi-domain SFT, the sequential training strategy of specific datasets in multiple phases can easily lead to significant performance drop of prior knowledge, which is well-known as Catastrophic Forgetting (Kaushik et al., 2021; McCloskey and Cohen, 1989), as illustrated in Table 1 and Figure 2. This problem hinders the versatile fine-tuning performance of a model across multiple domains (De Lange et al., 2021; Dong et al., 2023; Yuan et al., 2022). Therefore, mixing data from different domains is crucial for mitigating catastrophic forgetting during the SFT stage, enhancing the overall performance and adaptability.
@@
C2: Low Efficiency. Existing research on data composition during the supervised fine-tuning phase for LLMs is still in its initial stages, with most strategies based on heuristic or manually determined rules (Albalak et al., 2024; Dubey et al., 2024; Wang et al., 2023). One of the common baselines is defining domain weights referring to the natural domain sizes, which weights all the individual data points equally. Such approaches struggle to optimally balance different domains, failing to maximize the overall training effectiveness for multiple abilities. There lacks a well-defined methodology that efficiently enhances the versatile capabilities of LLMs across multiple domains during the SFT stage.
@@
C3: Low Flexibility in Domain Expansion. Existing SFT approaches for specific domain abilities typically pre-determine the proportions of different datasets according to prior experience (Azerbayev et al., 2023; Roziere et al., 2023). Such strategies lack the flexibility to dynamically adjust the data mixing ratios of different domains during the training process, which does not allow for real-time feedback from the LLMs in the current phase to inform and optimize the data composition. This static approach hinders the minimization of performance loss in other domains as LLMs undergo specialized training.
@@
To address these challenges, we introduce VersaTune, a novel data composition framework designed for enhancing models’ overall performances across different vertical domains during supervised fine-tuning. We first detect the proportion distribution of domain knowledge within the target model (Section 3.1), followed by data composition based on the existing distribution for multi-ability enhancement (Section 3.2.2) as well as flexible domain expansion (Section 3.2.3). VersaTune exhibits the following properties:
Efficient. VersaTune employs distribution consistency training of the domain knowledge proportion during models’ fine-tuning stage, providing an efficient data composition strategy for enhancing versatile capabilities (for C2).
Flexible. VersaTune can be flexibly adapted to scenarios that expand performance on specific domain tasks while minimizing the degradation of the model’s capabilities in non-target domains (for C1, C3).
Robust. The data composition strategy demonstrates significant improvements after fine-tuning in open-source models with parameter sizes ranging from 7B-14B, adding to the effectiveness of VersaTune (for C1, C2 and C3).
In summary, our contributions are as follows:
Challenges in Multi-Capability Training. Through a quantitative analysis of the performance losses in non-target domains resulting from existing training methods, we emphasize the challenge of data composition in multi-domain task training (Section 2).
@@
Multi-Capability Data Composition Framework. We introduce VersaTune, a data composition framework that leverages the intrinsic domain knowledge distribution of the model to proportion training data. VersaTune is designed to enhance overall performance across multiple domains while also providing flexibility for targeted improvements in specific domains (Section 3).
@@
Performance and Effectiveness. Our comprehensive evaluations across domains demonstrate that VersaTune can achieve an improvement of in versatile fine-tuning for multiple domains. Furthermore, when focusing on specific-domain fine-tuning, VersaTune maintains the training effectiveness in the target domain while reducing the performance degradation in other non-target domains by (Section 5).
2. Preliminary
2.1. Transformer Models
Transformers were initially introduced to address sequence modeling and transduction challenges, including language modeling and machine translation (Vaswani, 2017). The core of each Transformer layer consists of self-attention mechanisms and point-wise feed-forward networks. Self-attention network enables the model to assess the importance of various words in a sequence regardless of their positions (Zhang et al., 2019). It calculates a score for each word, indicating the degree of attention to be directed towards other sequence parts during encoding. This process involves query, key, and value weight sets that are optimized through training. Feed-forward network (FFN) applies the same linear transformation with a ReLU activation to each position individually. The FFN’s role is to independently transform each position’s representation, introducing non-linearity into the model and allowing it to learn complex functions (Miao et al., 2022). This component complements the self-attention mechanism by processing each input position in isolation, thus enhancing the model’s adaptability to diverse data types and tasks. These operations, predominantly dense algebraic computations such as matrix multiplications, lead to significant computational expenses and memory demands.
Domain of Training Dataset | Training Step | Variations in Comprehensive Domains () | Sum. () | |||||

### v5

@@ Abstract
As demonstrated by the proprietary Large Language Models (LLMs) such as GPT and Claude series, LLMs have the potential to achieve remarkable proficiency across a wide range of domains, including law, medicine, finance, science, code, etc., all within a single model. These capabilities are further augmented during the Supervised Fine-Tuning (SFT) phase. Despite their potential, existing work mainly focuses on domain-specific enhancements during fine-tuning, the challenge of which lies in catastrophic forgetting of knowledge across other domains. In this study, we introduce VersaTune, a novel data composition framework designed for enhancing LLMs’ overall multi-domain capabilities during training. We begin with detecting the distribution of domain-specific knowledge within the base model, followed by the training data composition that aligns with the model’s existing knowledge distribution. During the subsequent training process, domain weights are dynamically adjusted based on their learnable potential and forgetting degree. Experimental results indicate that VersaTune is effective in multi-domain fostering, with an improvement of 35.21% in the overall multi-ability performances compared to uniform domain weights. Furthermore, we find that Qwen-2.5-32B + VersaTune even surpasses frontier models, including GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 0.86%, 4.76% and 4.60%. Additionally, in scenarios where flexible expansion of a specific domain is required, VersaTune reduces the performance degradation in other domains by 38.77%, while preserving the training efficacy of the target domain.
VersaTune: An Efficient Data Composition Framework for Training Multi-Capability LLMs
Keer Lu†, Keshi Zhao†, Zhuoran Zhang‡, Zheng Liang§, Da Pan§, Shusen Zhang§, Xin Wu§, Guosheng Dong§, Bin Cui‡, Tengjiao Wang‡, Wentao Zhang† †Center for Data Science, Academy for Advanced Interdisciplinary Studies, Peking University, ‡School of CS & Key Lab of High Confidence Software Technologies (MOE), Peking University, §Baichuan Inc. {keer.lu, zhaoks}@stu.pku.edu.cn, {zhangzhuoran, bin.cui, tjwang, wentao.zhang}@pku.edu.cn {liangzheng, panda, zhangshusen, wuxin, dongguosheng}@baichuan-inc.com
1 Introduction
@@
Large Language Models (LLMs) have become a cornerstone in Artificial Intelligence (AI) Achiam et al. (2023); Dwivedi et al. (2021); Lewkowycz et al. (2022), particularly for Natural Language Processing tasks Brown et al. (2020); Devlin (2018); Radford et al. (2019), reshaping AI research and applications in domains such as law Cui et al. (2023), medicine Singhal et al. (2023); Thirunavukarasu et al. (2023), finance Li et al. (2023b); Wu et al. (2023), science Beltagy et al. (2019); Taylor et al. (2022) and code Liu et al. (2024b); Roziere et al. (2023). The success of LLMs stems from their capabilities to automatically learn and distill hierarchical data representations, making them highly effective for complex tasks Nie et al. (2023). In order to further enhance such abilities across these areas, LLMs typically undergo the supervised fine-tuning (SFT) stages on domain-specific datasets.
@@
As demonstrated by the robust performances of state-of-the-art LLMs such as GPT-4 Achiam et al. (2023) and Gemini Team et al. (2023), LLMs have the potential to master multiple tasks across all specific domains within a single model. However, most existing research on supervised fine-tuning tends to merely concentrate on a single ability of LLMs Dong et al. (2023); Xu et al. (2024), with the multi-domain performance on composite data of essentially different downstream tasks being less studied. We try to enhance the overall multitasking performance of LLMs across various domains by optimizing data mixing ratios during training:
@@
How to design a data composition strategy during SFT stages that could achieve overall multi-domain capabilities?
@@
Through analysis, we identified that the challenges associated with data composition strategies stem from the following three key aspects:
@@
C1: Catastrophic Forgetting. Given the fundamental differences between tasks of various domains, for multi-domain SFT, the sequential training strategy across multiple phases, where each phase exclusively utilizes a single-domain dataset for training, can easily lead to significant performance drop of prior knowledge, which is well-known as Catastrophic Forgetting Kaushik et al. (2021); McCloskey and Cohen (1989), as depicted in Table 2 and Figure 6. It hinders the versatile fine-tuning performance of a model across multiple domains De Lange et al. (2021); Dong et al. (2023); Yuan et al. (2022). Therefore, mixing data from different domains is crucial for mitigating catastrophic forgetting during training, enhancing the overall performance and adaptability.
@@
C2: Low Efficiency. Existing data composition research during the supervised fine-tuning phase for LLMs is still in its initial stages, with most strategies based on heuristic or manually determined rules Wang et al. (2023); Albalak et al. (2024); Dubey et al. (2024). One of the common baselines is defining domain weights referring to natural domain sizes, which weights all individual data points equally. Such approaches struggle to optimally balance different domains, failing to maximize the overall training effectiveness for multiple abilities. There lacks a well-defined methodology that efficiently enhances the versatile capabilities of LLMs across multiple domains during the SFT stage.
@@
C3: Low Flexibility in Domain Expansion. Existing SFT approaches for specific domain abilities typically pre-determine the proportions of different datasets according to prior experience Azerbayev et al. (2023); Roziere et al. (2023). Such strategies lack the flexibility to dynamically adjust the data mixing ratios of different domains during the training process, which does not allow for real-time feedback from LLMs to inform and optimize the data composition. This static approach hinders the minimization of performance loss in other domains as LLMs undergo specialized training.
@@
To address these challenges, we introduce VersaTune, a novel data composition framework to enhance models’ overall performances across different domains during supervised fine-tuning. We first detect the proportion distribution of domain knowledge within the target model (Section 2.1), followed by data composition based on the existing distribution for multi-ability enhancement (Section 2.2.2) and flexible domain expansion (Section 2.2.3). Our contributions are as follows:
Knowledge Consistency Training. We introduce the concept of knowledge consistency training for LLMs’ multi-capability development, which enables the model to continue learning from datasets that possess a knowledge distribution aligned with its pre-existing knowledge feature.
@@
Multi-Capability Data Composition Framework. We propose VersaTune, a novel data composition framework that leverages the model’s intrinsic domain knowledge distribution to optimize the training data proportion. VersaTune is designed to enhance the overall performance across multiple domains (Section 2.2.2), as well as to provide flexible expansion for specific domains while minimizing the performance degradation in other domains (Section 2.2.3).
@@
Performance and Effectiveness. Our extensive evaluations across domains demonstrate that VersaTune can achieve an improvement of 35.21% in versatile fine-tuning for multiple domains. Notably, we find that our Qwen-2.5-32B + VersaTune even outperforms frontier models including GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 0.86%, 4.76% and 4.60%. Furthermore, when focusing on specific-domain expansion, VersaTune maintains training effectiveness in the target domain while reducing performance degradation in other non-target domains by 38.77%.
2 VersaTune
In this section, we introduce VersaTune, a data composition framework designed for multi-capability training, aiming to effectively compose data from multiple domains and optimize the data proportion during training. Figure 1 presents the workflow of VersaTune, which generally contains two phases.
2.1 Phase 1: Domain Knowledge Detection
2.1.1 Knowledge Consistency Training
Previous research on data mixing ratios during the SFT phase for LLMs has predominantly focused on enhancing capabilities within a specific domain, often utilizing only data from that domain or employing heuristic, experience-based data proportions. We argue that such strategies can significantly impair the LLM’s abilities in other domains. In the fine-tuning stage, maintaining a robust overall capability across various domains is crucial.
What data mixing strategy effectively boosts the versatile performance of LLMs across domains during the SFT phase? We propose the statement:
Statement 1 An LLM fine-tuned with domain-specific data proportions that align with its pretrained output distributions will exhibit enhanced and balanced performance across these domains, compared to a model fine-tuned with a non-matching data distribution. Formally, the relationship can be represented as:
where denotes the set of all possible data points.
2.1.2 Knowledge Distribution Detection
Drawing on prior research into knowledge identification methods Gekhman et al. (2024); Zhao et al. (2023b) and training data inference strategies for LLMs Ding et al. (2022); Hayase et al. (2024), we propose a structured approach to efficiently detect domain knowledge based on statistics. The method involves the generation of textual outputs from the base model waiting for fine-tuning, followed by classification into predefined domains referring to a proprietary model . The process is repeated multiple times to ensure statistical robustness.
Assuming the data corpus contains distinct domains, as shown in Algorithm 1, we first prompt the base model poised for fine-tuning with the Beginning of Sequence (<BOS>) tokens to generate a set of data entries (Line 3). Subsequently, we employ a proprietary model to infer probabilities that these entries belong to each domain (Line 5-7). We then calculate a weighted average of the probability distributions for all data across these domains, thereby deriving the domain knowledge distribution of the current base model (Line 9). To ensure statistical robustness, the process is iteratively conducted times, and we use the mean of these iterations as the estimated result for knowledge distribution.
2.2 Phase 2: Fine-Tuning Multi-Ability LLMs Efficiently
Setting. Our goal is to construct a composite dataset covering specific domains, which can be denoted as with each tuple representing a specific domain and its corresponding proportion, such that training a model on dataset could achieve overall lower loss on a uniformly distributed composite target validation dataset or meet the flexible domain expansion while preserving the performances in other domains. The specialized capabilities of LLMs are measured using downstream tasks related to different domains (e.g., FinBen Xie et al. (2024a) for financial performances).
Input: Base model , Proprietary model , Hyperparameters: sample number , maximum iterations
Parameter: Data samples generated from
Output: Domain distribution
Define : domain probability distribution of data sample
2.2.1 Preliminary: Learnable Potential and Forgetting Degree of Knowledge
2.2.2 Effective Multi-Ability Fostering
Input: Base model to be fine-tuned , Domain reference loss , Hyperparameters: adjustment magnitude , training step number
Parameter: Data proportion of dataset
Output: Fine-tuned multi-ability model
Define : learnable potential of the current domain
As detailed in Algorithm 2, in the pursuit of balanced domain expertise enhancement, we first blended the domain proportions in accordance with the base model’s intrinsic domain knowledge distribution detected by Algorithm 1 (Line 1). Then at each training step , we assigned a learnable potential score to each domain based on the methodology outlined in Equation 2. These scores were then utilized to fine-tune the representation of each domain within the composite SFT dataset, ensuring a balanced development of competencies across all domains throughout the training process (Line 3-7). At the same time, the parameters of model are updated based on the gradients computed through backpropagation (Line 8). This adaptive approach is imperative to harmonize the progression of capabilities in different domains and to optimize the model’s performance on multiple tasks.
2.2.3 Flexible Domain Expansion
When conducting fine-tuning on a pretrained model, there are instances where we aim to particularly enhance models’ performance on specific domain tasks. Consequently, our algorithmic framework ought to possess the flexibility to accommodate domain expansion and generalize effectively. Building upon Statement 1, we present the following statement tailored for domain expansion:
Statement 2
where is the data distribution in the domain , and is the adjustment factor.
where is the increased adjustment factor, and is the original ratio of other domain knowledge relative to . Algorithm implementation details and hyper-parameter settings are provided in Section B.1 and Section C.2.1.
3 Experiments and Results
In this section, we describe details of our experimental setup (Section 3.1), the baseline methods we use for comparison (Section 3.2), and experimental results (Section 3.3).
Frontier Models | |||||||||||||
GPT-4o | – | 79.00 | 57.41 | 81.92 | 74.60 | 64.58 | 66.25 | 72.54 | 85.47 | 88.40 | 75.50 | 71.82 | 90.56 |
Claude3.5-Sonnet | – | 77.60 | 40.73 | 76.38 | 68.80 | 65.90 | 62.55 | 68.72 | 84.24 | 84.07 | 80.48 | 75.63 | 89.12 |
DeepSeek-V3 | – | 65.46 | 52.25 | 78.82 | 74.30 | 68.15 | 75.03 | 69.58 | 82.90 | 65.20 | 75.40 | 79.60 | 88.90 |
Open-Sourced Base Models | |||||||||||||
LLaMA-2-7B | Uniform Distribution | 15.71 | 30.72 | 23.45 | 27.57 | 33.50 | 2.71 | 9.30 | 42.89 | 5.67 | 3.44 | 20.16 | 71.40 |
LLaMA-2-13B | Uniform Distribution | 47.66 | 34.85 | 32.98 | 36.54 | 37.54 | 32.85 | 45.72 | 50.77 | 36.54 | 38.55 | 36.89 | 73.50 |
VersaTune | 60.59↑ | 46.58↑ | 50.24↑ | 45.15↑ | 65.84↑ | 78.68↑ | 62.89↑ | 82.86↑ | 82.64↑ | 81.48↑ | 55.52↑ | 75.98↑ | |
Qwen-2.5-32B | Uniform Distribution | 68.86 | 45.28 | 72.34 | 68.18 | 68.03 | 75.14 | 58.30 | 80.17 | 78.59 | 71.04 | 75.26 | 84.40 |
Inverse Distribution | 62.93↓ | 42.05↓ | 68.80↓ | 66.09↓ | 66.80↓ | 73.93↓ | 52.94↓ | 79.31↓ | 74.44↓ | 70.71↓ | 75.00↓ | 83.80↓ | |
VersaTune | 75.67↑ | 56.76↑ | 78.72↑ | 72.36↑ | 70.50↑ | 78.80↑ | 70.77↑ | 85.23↑ | 86.60↑ | 79.89↑ | 75.81↑ | 84.75↑ |
3.1 Experimental Setup
Datasets. For training, we have collected datasets spanning 6 domains for supervised fine-tuning, including Lawyer-Instruct111https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-Instruct, the training portion of MedQA Jin et al. (2020), Finance Alpaca222https://huggingface.co/datasets/gbharti/finance-alpaca, Sonnet3.5 Science Conversations333https://huggingface.co/datasets/jeffmeloy/sonnet3.5_science_conversations, Code Alpaca444https://github.com/sahil280114/codealpaca and Alpaca Taori et al. (2023), denoted as , to represent SFT datasets with respect to law, medicine, finance, science, code as well as general capabilities. In order to prevent domain overlap, we curated the Alpaca dataset by excluding data pertaining to the other specific five domains, keeping only the general domain instances unrelated to them. More details can be found in Section C.2. For evaluation, we assess the model performances on downstream tasks across various domains, using two relevant benchmarks for each domain, with details provided in Section C.3.
Models and Implementation. We employ LLaMA Dubey et al. (2024); Touvron et al. (2023a, b) and Qwen Bai et al. (2023); Yang et al. (2024) series as our pretrained language models . During the fine-tuning procedure, we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 4 epochs. To maintain consistency, the total volume of training data across domains was controlled to 60,000 per epoch. We conducted all fine-tuning and evaluation experiments on NVIDIA RTX H800. Details of the experimental settings can be found in Appendix C.
3.2 Baselines
We compare VersaTune with the following baselines. For the scenario of effective multi-ability fostering: (1) The simplest baseline is uniform distribution, where each domain has an equal weight proportion. (2) Inverse distribution assigns the proportionate weights to each domain in an inverse manner to the detected knowledge distribution. (3) Frontier models contain GPT-4o Hurst et al. (2024), Claude3.5-Sonnet Anthropic (2024) and DeepSeek-V3 Liu et al. (2024a). Under the case of flexible domain expansion: (1) 100% specific domain strategy is a common practice to employ datasets consisting exclusively of data from a single domain during the fine-tuning stage. (2) Domain increase with uniform distribution of remainder elevates the proportion of a specific domain, while the remaining domains receive the balance of the distribution in an evenly distributed manner.
3.3 Results
VersaTune is efficient across different models in both scenarios. For the scenario of multi-capability fostering, Table 1 shows that VersaTune consistently outperforms other baseline methods across different models in terms of domain-specific capabilities. Compared to the uniform distribution of data across domains, VersaTune enhances downstream task performances by , which further underscores the effectiveness of our data composition strategy for enhancing the model’s overall multi-domain capabilities during the supervised fine-tuning phase. Moreover, Qwen-2.5-32B + VersaTune has the potential to surpass frontier models under medical scenarios, achieving average improvements over GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 0.86%, 4.76% and 4.60%. Since we have not conducted domain-specific refinement for domains outside the current five specific domains, the models’ performance gains on general benchmarks are not as noticeable. For domain expansion scenarios, VersaTune has nearly maintained training efficiency while reducing the model’s loss of competencies in other domains by comparing to 100% specific domain fine-tuning, as depicted in Table 7, where we averaged the experimental results from Qwen-2.5-7B and Qwen-2.5-14B. Detailed results and analysis can be found in Appendix D.
Knowledge consistency training boosts performance. In Table 1, we present the experimental results of data composition strategies that allocate domain data in a manner inversely proportional to the pre-existing knowledge distribution detected within each domain. As expected, the inverse distribution strategy yielded even lower performance compared to the simplest approach of uniform distribution, which evenly distributes data across all domains. This finding underscores the importance of aligning domain data ratios with the inherent knowledge distribution of the model during training, which proves the efficacy of knowledge consistency training stated in Section 2.1.1.
4 Ablations and Analysis
Previously in Section 3, we have demonstrated the effectiveness of VersaTune in enhancing multiple abilities and enabling flexible domain expansion during the SFT phase. In this section, we perform an in-depth analysis of VersaTune, where we ablate the components of (1) dynamic adaptation in Algorithm 2, and (2) the criteria for determining the upper limit of domain expansion in Algorithm 3.
Dynamic adjustment enhances the robustness. During the process of cultivating multiple capabilities, we compared VersaTune with fixed domain weights referring to the knowledge distribution obtained from probing the target model prior to supervised fine-tuning, namely VersaTune Constant, to ablate the component of dynamic adaptation in Algorithm 2. Table 5, Figure 2, and Figure 8 demonstrated the high robustness of VersaTune, which dynamically adjusts domain weights throughout the training process by continuously monitoring the learnable potential within each domain. In contrast, training with fixed domain weights exhibits certain fluctuations. A key reason for this phenomenon is the distribution of domain knowledge mastered by the model changes during training, and the learning efficiency varies among domains. Therefore, dynamically adjusting domain data weights based on the model’s feedback at different stages of training is crucial. More experimental results can be found in Section D.1.
Establishing proportion thresholds for specific domains counts during domain expansion. We consider conducting a comparative analysis between the outcomes of VersaTune and those implementing an unconditional dynamic increase of the specific domain, where we remove the implementation of Line 8 in Algorithm 3. Figure 4 shows that the criteria for determining the upper limit on the proportion of a specific domain during domain expansion, has mitigated the loss of capabilities in other domains experienced by the target model during the fine-tuning process. Concurrently, it ensures gains in the capacity for the current domain of interest. We speculate that it is because by the later stages of fine-tuning, models’ proficiency in the target domain approaches saturation. Further increasing the proportion of the current domain provides diminishing returns and can lead to a significant loss of performance in other domains. Detailed analysis are provided in Section D.2.2.
5 Related Work
Data Reweighting for LLM Training. Data reweighting maintains full access to the entire dataset while adjusts the relative importance of each instance for various target tasks, which is essential for both pretraining and fine-tuning stages of LLMs Wang et al. (2023). During the pretraining stage, DoReMi Xie et al. (2024b) and DoGE Fan et al. (2023) employ lightweight proxy models to estimate weights for different data sources, which are subsequently applied to the formal training of LLMs. Furthermore, Sheared LLaMA Xia et al. (2023) implements an online variant of DoReMi. As for the SFT phase, Dong et al. Dong et al. (2023) focus on enhancing the model’s math reasoning, coding, and human-aligning abilities through a dual-stage mixed fine-tuning strategy. However, the mixing ratios for different domains rely heavily on empirical methods, and the covered domains are not holistic. We provide a comprehensive overview of the model’s capabilities across domains during the SFT stage and proposes appropriate and holistic multi-ability fine-tuning methods.
Knowledge Detection in LLMs. Investigating the knowledge contained in LLMs is essential for guiding their subsequent training Chang et al. (2024). The knowledge encompasses multiple dimensions, such as different domain sources and task attributes. Existing work on LLM knowledge detection primarily focuses on prompting and calibration. Directly prompting the model to generate sequences and extracting confidence scores from the model Gekhman et al. (2024); Kadavath et al. (2022); Kuhn et al. (2023); Manakul et al. (2023) is a common strategy. However, such approaches highly depend on prompt design and task selection, introducing bias into the assessment. Other studies have attempted to infer the training data mixtures used in previous training stages Antoniades et al. (2024); Hayase et al. (2024); Hu et al. (2022); Ye et al. (2022). The essence of these studies is to evaluate the current knowledge state of the models and provide targeted strategies for data organization and management in subsequent training phases.
6 Conclusion
In this work, we introduce VersaTune, a noval data composition framework designed to enhance the multi-domain capabilities of models during the supervised fine-tuning phase of LLMs, which is based on the domain knowledge distribution of the target model. Experimental results have demonstrated that VersaTune achieves excellent training outcomes in both scenarios of overall multi-domain enhancement and flexible domain expansion.
7 Limitations
There are some limitations in our work. Firstly, the classification framework of vertical domains may not be comprehensive in scope, and since the classifier relies on advanced language models, it cannot guarantee absolute accuracy in classification. Additionally, when computing the learnable potential and forgetting degree of knowledge, to balance the computational cost and effectiveness, we employ a lightweight proxy model to for calculation, yet it does not fully represent the performance tendencies of the target model during actual evaluating.
8 Ethical Considerations
Integrating Large Language Models (LLMs) into domain reweighting settings holds potential for improving multi-domain capabilities of models, while it also brings several ethical considerations that must be addressed to ensure responsible and beneficial use. VersaTune dynamically adjusts data distribution based on the model’s existing knowledge to ensure fairness and avoid biases that could arise from the data composition process. Additionally, VersaTune adhere to privacy standards by merely utilizing open-sourced datasets, ensuring that personal data used in the training process is anonymized and securely handled to protect individual privacy rights.
Appendix
- 1 Introduction
- 2 VersaTune
- 3 Experiments and Results
- 4 Ablations and Analysis
- 5 Related Work
- 6 Conclusion
- 7 Limitations
- 8 Ethical Considerations
- A Background and Discussion
- B Method Details
- C Experiments Details
- D More Experiment Results
- E Prompts
Appendix A Background and Discussion
In this section, we provide the background information and design motivation for our VersaTune.
A.1 Pretraining and Supervised Fine-Tuning
The training process of Large Language Models (LLMs) generally involves the pretraining and fine-tuning stages. We have outlined several concepts about LLMs training.
A.1.1 Pretraining
Large Language Models (LLMs) establish basic knowledge abilities, including language understanding and text generation, during the pretraining stage Brown et al. (2020). In this stage, LLMs engage in unsupervised training through the processing of extensive raw text corpora, thereby enhancing their capabilities in language modeling. For a given sequence , the typical task for LLMs involves the prediction of the subsequent token given the preceding tokens as contextual input. The goal is to maximize the likelihood function presented in Equation 6:
Beginning of Sequence (<BOS>) During the above process, the Beginning of Sequence (<BOS>) token plays an important role, which serves as a signal to the model that the input sequence is starting Brown et al. (2020); Li et al. (2023a); Wu et al. (2025). It can be thought of as a special marker that indicates the start of a new sequence, allowing the model to reset its context and begin processing a new piece of text. In the context of pretraining, the <BOS> token is used to initialize the input to the model, and it can be concatenated with the actual text data to form the input sequence. This token helps the model to differentiate between the start of a new input and the continuation of an existing one. It is particularly crucial in tasks where the model needs to generate text or understand the beginning of a new sentence or document, which helps the model to learn the boundaries of text sequences and to better model the statistical properties of the language data it is trained on. The use of <BOS> tokens, along with other special tokens like <EOS> (End of Sequence), helps the model to learn the boundaries of text sequences and to better model the statistical properties of the language data it is trained on.
A.1.2 Supervised Fine-Tuning
The Supervised Fine-Tuning (SFT) stage of a Large Language Model (LLM) involves further training to refine the model’s task-solving capabilities and ensure greater alignment with human instructions Zhao et al. (2023a). While recent research has delved into exploring fine-tuning methods for multi-task enhancement Dong et al. (2023); Sanh et al. (2021), they are still in their early stages. However, as shown by proprietary models such as GPT-4 Achiam et al. (2023), Gemini Team et al. (2023), and DeepSeek series Liu et al. (2024a), which exhibit outstanding multi-task performance, improving a model’s versatile capabilities across various domains during the SFT phase is crucial. Therefore, our work systematically investigates methods to enhance multi-domain performance during the SFT stage to bridge this gap.
A.2 Analysis on Catastrophic Forgetting
Target Domain | Training Step | Variations in Comprehensive Domains () | Sum. () | |||||